At present time a man is four times as old as his son. Six years ago he was as 10 times as old. Find their present age.
Answer by stanbon(75887)

At present time a man is four times as old as his son. Six years ago he was as 10 times as old. Find their present age.
--------------
Equations:
m = 4s
m-6 = 10(s-6)
----
Substitute for "m" and solve for "s":
4s-6 = 10s-60
6s = 54
s = 9 yrs (son's age now)
m = 4s = 36 (father's age now)
------------
Cheers,
Stan H.
-----------------